Find the indicated IQ score. The graph to the right depicts IQ scores of​ adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15. with the distribution of 0.55
1
Expert's answer
2020-04-13T14:47:10-0400
"P(X>x)=0.55"

The z-score of a normally distributed random variable, X, with mean μ and standard deviation σ, at the point x, is given as:


"Z={X-\mu \over \sigma}"

"P(X\leq x)=1-P(X>x)=1-0.55=0.45"

"P(Z\leq {x-\mu \over \sigma})=0.45=>{x-\mu \over \sigma}\approx-0.12566"

"x\approx15(-0.12566)+100\approx98.1151"

"x\approx98.1"